A man has a 16 quart radiator that now contains 16 quarts of 20 percent antifreeze. How much of the current mixture has to be removed and replaced with pure antifreeze to raise the level of concentration to 35 percent?
Answer by ankor@dixie-net.com(22740)

A man has a 16 quart radiator that now contains 16 quarts of 20 percent antifreeze.
How much of the current mixture has to be removed and replaced with pure antifreeze to raise the level of concentration to 35 percent?
:
let x = amt removed, and amt of pure antifreeze to replace it
:
.20(16-x) + x = .35(16)
3.2 - .2x + x = 5.6
.8x = 5.6 - 3.2
.8x = 2.4
x = 2.5/.8
x = 3 qts removed and replaced with 3 qts of pure antifreeze